# Uncertainty-driven Adaptive Exploration

### v3

@@ 3 Experimental Evaluation
Finally, for interest, we also create an agent that uses Noisy Nets [fortunato2018noisy] to perturb the actor’s network and achieve more efficient exploration. We compare against this TD3+Noisy Nets agent, since Noisy Nets is a popular choice for exploring robotic domains. Note that Noisy Nets does not utilize any definition of uncertainty—it is just an “improved” TD3 agent in terms of exploration.
We conduct our experiments in standard MuJoCo domains [towers2024gymnasium]. Specifically, we used the five domains considered the most difficult according to the official documentation. [^10] https://gymnasium.farama.org/environments/mujoco/ . We did not use [gymnasium_robotics2023github] since its domains do not have the “early termination” condition. Such domains do not satisfy Definition 2.1. However, the ADEU agents still achieve (as later described) superior performance against their opponents. In these domains, the agent should explore the continuous multidimensional action space to find the best policy. A dense reward function is used to describe these robotic tasks. However, ADEU can also be used in environments with a sparse reward function, when using appropriate uncertainty-measuring mechanisms e.g., [Badia2020NeverGiveUp, nikolov2018informationdirected].
@@
Results in the DeepSea domain
@@
Figure 4: Timesteps required for each algorithm to solve the DeepSea domain for $N\times N$ grid. The lower the curve, the better the performance of the algorithm.
@@
From the conducted experiments and the corresponding results, it is clear that the ADEU framework, paired with efficient uncertainty-measuring mechanisms, can be utilized for both demanding control tasks and domains requiring deep exploration.
4 Related Work
@@
Intrinsic Motivation: Several IM approaches have been tested in robotic domains.
For instance, [Lobel2023FlippingCoins] design an efficient intrinsic motivation function that can be utilized as an uncertainty mechanism in ADEU. They test their method in a sparse reward variance of the Fetch domain; an easier domain than standard robotics. Then, [mahankali2024random] adds a random reward to the sparse reward function, to guide the agent to explore different parts of the environment. This can be used by ADEU in extremely sparse reward function domains.
Epistemic Uncertainty: As mentioned earlier, UCB [Chen2017ucb, lee2021sunrise] is commonly used in continuous action-space robotic domains.
We also created an instance of ADEU using this ensemble mechanism to track uncertainty (the ADEU instance with $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$ in Sec. 3) and test it against a UCB implementation built on top of a TD3 agent. da2020uncertainty “asks” a mentor for guidance when the uncertainty is high; uncertainty was measured in a similar way.
Active Exploration: Several works [russo2023active1, hazan2019active2, hong2020active4, nikolov2018informationdirected] have proposed methods to efficiently explore the environment to derive an $\epsilon$-optimal policy, while also obtaining a low sample complexity. Some methods [russo2023active1, hazan2019active2], where tested in tabular or “discretized” environments. Hence, these works tackle a distinct exploration problem, and there is no straightforward comparison with ADEU. Other works [hong2020active4, nikolov2018informationdirected] incorporate these ideas in the classic DRL testbeds either by proposing mechanisms to motivate efficient exploration [nikolov2018informationdirected] or by changing the reward function that the agent aims to maximize [hong2020active4]. Both mechanisms can be used in the ADEU framework.
@@
5 Conclusions and Future Work
In this paper, we put forward a generic adaptive exploration framework that uses a generic uncertainty-based action selection mechanism to decide in a principled manner when to alternate between exploration and exploitation. The generality of that mechanism allows (i) our framework to effectively incorporate any uncertainty-measuring mechanism of choice; and (ii) existing adaptive exploration techniques to be viewed as special cases of our framework. In problems requiring adaptive exploration, ADEU is a promising and easily deployable solution. Additionally, it allows the user to select an existing uncertainty measurement mechanism or to define a heuristic one tailored to its problem.Our experiments verified the superiority of ADEU in difficult robotic testbeds.
Ongoing and future work includes extending ADEU in various directions. To begin, we intend to perform tests to further verify ADEU’s ability to exploit background knowledge; and its ability to recover from being fed with incorrect background information or to detect and escape clear suboptimal trajectories by optimizing the rollout procedure. In addition, to make the learning procedure more robust, we will modify the learning process by weighting the learned experience before backpropagating it in the network, similar to [lee2021sunrise]. As mentioned earlier, we did not modify the learning procedure in this work, since ADEU does not exhibit instabilities, and we also aim to keep this work “exploration-only”. Moreover, we are actively working to extend ADEU to guarantee safe exploration [wachi2023safe, ijcai2022p3o]. In ongoing work, we are calculating the upper bound of allowed exploration in neighboring safe states, so that ADEU can guarantee safe exploration. We also aim to extend ADEU to multiagent environments. By leveraging agents’ independent exploration, ADEU agents can adaptively explore the environment. Furthermore, by allowing a centralized agent to define $f_{i}(s)$ for each agent $i$, we can reduce the non-stationarity issues that usually arise in multiagent RL. Finally, we intend to design mechanisms to appropriately set the third and fourth moments of $D$, to guide exploration to particular directions.
Acknowledgments and Disclosure of Funding
The research described in this paper was carried out within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union - NextGenerationEU (Implementation Body: HFRI. Project name: DEEP-REBAYES. HFRI Project Number 15430).
@@ Appendix A Technical details regarding the “modified Frozen Lake” motivating example
Figure 5: Environment used in our motivation example. In this particular figure, the dimensions of the grid are $10\times 10$. In our example, we use a $1500\times 1500$ grid.
@@
In the case of a “hole”, the agent receives a penalty (=10)(=-10), and the episode is terminated; while when the agent reaches the goal terminal state, it receives a high reward—equal to $10^{5}$ in our case.
The theoretical maximum reward in this environment when $N=1500$ is calculated as the sum of the rewards in the states of the sole trajectory i.e., states with: idid 0 and idsids between $1$, and (N2)N+2(N-2)N+2 with step $N$ and idsids between (N2)N+2(N-2)N+2 and (N2)N+N(N-2)N+N and state $(N^{2})-1$ ($2999$ in total). The evaluation reward for those states is $104,489$ (unsmoothed—$4,489$ for the trajectory + $10^{5}$ for finding the treasure) or $104,487$ (smoothed), equal to the one obtained by ADEU (see Table 3).
As an intrinsic reward mechanism for the IM agent, we use the visitation frequency, calculated as $\frac{\beta}{\sqrt{n(s)}}$, where is the exact visitation count $n(s)$. This approach is similar to [tang2017exploration], but without the need to approximate the visitation frequency, as in this tabular scenario we can in fact keep track of the times that an agent has visited a state $s$. As such, the reward function used for the IM agent is :
@@
Figure 6: The “shifted” sigmoid function compared to the classic sigmoid.
@@ Appendix B Ablation Study on the Normalizer Function
Figure 7: The two sigmoids tested in our main experiments
While this approach has desirable properties, it may lead to unintended consequences. Specifically, when uncertainty is low, the agent may repeatedly generate identical or highly similar experiences. This lack of diversity in the training data can increase the risk of catastrophic forgetting, since the agent produces correlated experience [Bruin2015catastrophic_forgetting_2]—at least for the trusted policy.
@@ Appendix C Implementation details on RND’s networks
Appendix D Hyperparameters and Hardware details
@@
Table 6: Parameters used in our experiments
@@
All of our code was implemented in Python using the $3.11$ version. In addition, we use the $2.3.1$ version of pyTorch [paszke2019pytorch] and the $1.26.1$ version of numpy [harris2020numpy]. We also use the $1.1$ version of gymnasium [towers2024gymnasium], and we ran our experiments using the latest ($v5$) version of the domains.
Appendix E Extended results
@@
Figure 8: The cumulative episodic rewards yielded by each exploration strategy/framework in the tested domains. Results are averaged over $15$ seeds. Shaded areas correspond to the standard deviation.
Appendix F Recreating ez-greedy as an ADEU instance
As already discussed, a Distribution $D$, a policy $\pi(s)$, and an uncertainty-measuring mechanism $f$ are required, to define any instance of an ADEU agent. To emulate dabney2021ez_greedy, we will define the following parameters:
Table 7: Emulating $\epsilon z$-greedy
@@ Appendix F Recreating ez-greedy as an ADEU instance
where $f(s,n,\cdot)$ is the uncertainty-measuring mechanism described in Algorithm 1, $Q(s,a_{i})$ is the discrete action space Q-Function for each action $i$ derived from any DRL algorithm of choice, and $U(A)$ is the mechanism that calculates the options as described in [dabney2021ez_greedy].
@@
1:Inputs: $\epsilon$ (float): probability to begin an option (see [dabney2021ez_greedy]), $n$ (integer $\geq 0$): steps exploring using an option, explored (boolean): indicates whether the agent explored or not during the previous step.
@@
To be consistent with the previous notation, in that case $f()$ is called with n=n=previous returned $n$, $\epsilon=\rho$, and explored = True if the agent sampled $U(A)$, otherwise it is set to False. For the rest “implementation details” please refer to the corresponding Section and dabney2021ez_greedy.

### v4

@@ 3. Experimental Evaluation
Finally, for interest, we also create an agent that uses Noisy Nets fortunato2018noisy to perturb the actor’s network and achieve more efficient exploration. We compare against this TD3+Noisy Nets agent, since Noisy Nets is a popular choice for exploring robotic domains. Note that Noisy Nets does not utilize any definition of uncertainty—it is just an “improved” TD3 agent in terms of exploration.
We conduct our experiments in standard MuJoCo domains towers2024gymnasium. Specifically, we used the five domains considered the most difficult according to the official documentation. [^10] https://gymnasium.farama.org/environments/mujoco/ . We did not use gymnasium_robotics2023github since its domains do not have the “early termination” condition. Such domains do not satisfy Definition 1. However, the ADEU agents still achieve (as later described) superior performance against their opponents. In these domains, the agent should explore the continuous multidimensional action space to find the best policy. A dense reward function is used to describe these robotic tasks. However, ADEU can also be used in environments with a sparse reward function, when using appropriate uncertainty-measuring mechanisms e.g., Badia2020NeverGiveUp; nikolov2018informationdirected.
@@
Results in the DeepSea domain:
@@
Figure 4. Timesteps required for each algorithm to solve the DeepSea domain for $N\times N$ grid. The lower the curve, the better the performance of the algorithm.
@@
To conclude the experimental section, it is clear that the ADEU framework, paired with efficient uncertainty-measuring mechanisms, can be utilized for both demanding control tasks and domains requiring deep exploration.
4. Related Work
@@
Intrinsic Motivation: Several IM approaches have been tested in robotic domains. For instance, Lobel2023FlippingCoins design an efficient intrinsic motivation function that can be utilized as an uncertainty mechanism in ADEU. They test their method in a sparse reward variance of the Fetch domain; an easier domain than standard robotics. Then, mahankali2024random adds a random reward to the sparse reward function, to guide the agent to explore different parts of the environment. This can be used by ADEU in extremely sparse reward function domains.
Epistemic Uncertainty: As mentioned earlier, UCB Chen2017ucb; lee2021sunrise is commonly used in continuous action-space robotic domains. We also created an instance of ADEU using this ensemble mechanism to track uncertainty (the ADEU instance with $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$ in Sec. 3) and test it against a UCB implementation built on top of a TD3 agent. da2020uncertainty “asks” a mentor for guidance when the uncertainty is high; uncertainty was measured in a similar way.
Active Exploration: Several works russo2023active1; hazan2019active2; hong2020active4; nikolov2018informationdirected have proposed methods to efficiently explore the environment to derive an $\epsilon$-optimal policy, while also obtaining a low sample complexity. Some methods russo2023active1; hazan2019active2, where tested in tabular or “discretized” environments. Hence, these works tackle a distinct exploration problem, and there is no straightforward comparison with ADEU. Other works hong2020active4; nikolov2018informationdirected incorporate these ideas in the classic DRL testbeds either by proposing mechanisms to motivate efficient exploration nikolov2018informationdirected or by changing the reward function that the agent aims to maximize hong2020active4. Both mechanisms can be used in the ADEU framework.
@@
5. Conclusions and Future Work
In this paper, we put forward a generic adaptive exploration framework that uses a generic uncertainty-based action selection mechanism to decide in a principled manner when to alternate between exploration and exploitation. The generality of that mechanism allows (i) our framework to effectively incorporate any uncertainty-measuring mechanism of choice; and (ii) existing adaptive exploration techniques to be viewed as special cases of our framework. In problems requiring adaptive exploration, ADEU is a promising and easily deployable solution. Additionally, it allows the user to select an existing uncertainty measurement mechanism or to define a heuristic one tailored to its problem. Our experiments verified the superiority of ADEU in difficult testbeds.
Ongoing and future work includes extending ADEU in various directions. To begin, we intend to perform tests to further verify ADEU’s ability to exploit background knowledge; and its ability to recover from being fed with incorrect background information or to detect and escape clear suboptimal trajectories by optimizing the rollout procedure. Moreover, we are actively working to extend ADEU to guarantee safe exploration wachi2023safe; ijcai2022p3o. In ongoing work, we are calculating the upper bound of allowed exploration in neighboring safe states, so that ADEU can guarantee safe exploration. We also aim to extend ADEU to multiagent environments. By leveraging agents’ independent exploration, ADEU agents can adaptively explore the environment. Furthermore, by allowing a centralized agent to define $f_{i}(s)$ for each agent $i$, we can reduce the non-stationarity issues that usually arise in multiagent RL. Finally, we intend to design mechanisms to appropriately set the third and fourth moments of $D$, to guide exploration to particular directions.
@@ Appendix A Technical details regarding the “modified Frozen Lake” motivating example
Figure 5. Environment used in our motivation example. In this particular figure, the dimensions of the grid are $10\times 10$. In our example, we use a $1500\times 1500$ grid.
@@
In the case of a “hole”, the agent receives a penalty (=10)(=-10), and the episode is terminated; while when the agent reaches the goal terminal state, it receives a high reward—equal to $10^{4}$ in our case.
As an intrinsic reward mechanism for the IM agent, we use the visitation frequency, calculated as $\frac{\beta}{\sqrt{n(s)}}$, where is the exact visitation count $n(s)$. This approach is similar to tang2017exploration, but without the need to approximate the visitation frequency, as in this tabular scenario we can in fact keep track of the times that an agent has visited a state $s$. As such, the reward function used for the IM agent is :
@@
Figure 6. The “shifted” sigmoid function compared to the classic sigmoid.
@@ Appendix B Ablation Study on the Normalizer Function
Figure 7. The two sigmoids tested in our main experiments
While this approach has desirable properties, it may lead to unintended consequences. Specifically, when uncertainty is low, the agent may repeatedly generate identical or highly similar experiences. This lack of diversity in the training data can increase the risk of catastrophic forgetting, since the agent produces correlated experience Bruin2015catastrophic_forgetting_2—at least for the trusted policy.
@@ Appendix C Implementation details on RND’s networks
Appendix D Proof of Theorem 1
Proof.
To prove Theorem 4, we will use mathematical induction.
Appendix E Generalizing Definition 2.1:
Definition 0 (Non-Terminal-Hyper-State).
Definition 0 (Terminal-Hyper-States).
Using the aforementioned definitions of Hyper-States, Definition 1 can be generalized also to include “Sparse Increasing Reward functions Single Agent Games” as follows:
Definition 0 (Sparse Increasing Reward Functions Single Agent games).
We define as “Sparse Increasing Reward Function Single Agent Game” a game that replaces states with Hyper-States in Definition 1. In more detail, $s^{*}$ states are replaced by Nom-Terminal-Hyper-State with $r(s_{d},a_{d})>0$, $s^{!}$ by Non-Terminal-Hyper-State with $r(s_{d},a_{d})<0$ and all the other states are replaced by Terminal-Hyper-States.
Corollary 0.
Theorem 4 holds for finding the best hyper-action in Sparse Increasing Reward Functions Single Agent Games. A hyper-action can be defined as a sequence of actions, exactly like “options”. The uncertainty-measuring mechanism in ADEU variances that act in games satisfying Definition 7 should calculate uncertainty in Hyper-States (i.e., a batch of non-terminal states).
Appendix F Hyperparameters and Hardware details
@@
Table 6. Parameters used in our experiments
@@
All of our code was implemented in Python using the $3.11$ version. In addition, we use the $2.3.1$ version of pyTorch paszke2019pytorch and the $1.26.1$ version of numpy harris2020numpy. We also use the $1.1$ version of gymnasium towers2024gymnasium, and we ran our experiments using the latest ($v5$) version of the domains.
Appendix G Extended results
@@
Figure 8. The cumulative episodic rewards yielded by each exploration strategy/framework in the tested domains. Results are averaged over $15$ seeds. Shaded areas correspond to the standard deviation.
Appendix H Recreating ez-greedy as an ADEU instance
[…]
Table 7. Emulating $\epsilon z$-greedy
@@ Appendix H Recreating ez-greedy as an ADEU instance
where $f(s,n,\cdot)$ is the uncertainty-measuring mechanism described in Algorithm 1, $Q(s,a_{i})$ is the discrete action space Q-Function for each action $i$ derived from any DRL algorithm of choice, and $U(A)$ is the mechanism that calculates the options as described in dabney2021ez_greedy.
@@
1:Inputs: $\epsilon$ (float): probability to begin an option (see dabney2021ez_greedy), $n$ (integer $\geq 0$): steps exploring using an option, explored (boolean): indicates whether the agent explored or not during the previous step.
@@
To be consistent with the previous notation, in that case $f()$ is called with n=n=previous returned $n$, $\epsilon=\rho$, and explored = True if the agent sampled $U(A)$, otherwise it is set to False. For the rest “implementation details” please refer to dabney2021ez_greedy.